OpenStudy (anonymous):

can somebody solve 6(sqrtx+5)+1=55

OpenStudy (anonymous):

6sqrtx+30+1=55 6sqrtx=24 sqrtx = 4 x = 16

OpenStudy (anonymous):

(sqrtx+5) = 54/6 = 9 sqrtx = 4 x = 16